Find the additive inverse of each number
1. -2 - 3i
2. 1 + 4i
3. 5 - 3i
4. -7 + 2i

Answer by drk(1908)

Given a number, a, the additive inverse is -a.
The same should apply to a+bi. The additive inverse should be -(a+bi) = -a - bi
Our problems are:
1. -2 - 3i
additive inverse is -(-2 - 3i) = 2 + 3i
2. 1 + 4i
additive inverse is -1 - 4i
3. 5 - 3i
additive inverse is -5 + 3i
4. -7 + 2i
additive inverse is 7 - 2i